# Page curve entanglement dynamics in an analytically solvable model

### v2

@@ Abstract
One related interesting property of the state in the environment is that its energy variance decreases after the Page time $t_{\rm P}$ (Fig. 8), corresponding to fewer eigenstates of $H_{\rm env}$ contributing to it. Asypmotically for $t\gg t_{\rm P}$ the energy variance $\Delta H_{\rm env}^{2}$ becomes independent of the number of particles $M$ in the environment (35), which is different from the usual linear dependence on $M$ for independently injected particles into ${\cal H}_{\rm env}$. Such states with reduced energy uncertainty at nonzero excitation energy could be of interest experimentally.
@@
Acknowledgments. I thank R. Jha and L. Tagliacozzo for valuable discussions. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)-217133147/SFB 1073 (Project No. B07). It was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, and at the Kavli Institute for Theoretical Physics (KITP), supported by grants NSF PHY-1748958 and PHY-2309135, whose hospitality is gratefully acknowledged.

### v3

@@ Abstract
Finally, one related interesting property of the state in the environment is that its energy variance decreases after the Page time (Fig. 10), corresponding to fewer eigenstates of $H_{\rm env}$ contributing to it. Asymptotically for $t\gg t_{\rm P}$ the energy variance $\Delta H_{\rm env}^{2}$ becomes independent of the number of particles $M$ in the environment (65), which is different from the usual linear dependence on $M$ for independently injected particles into ${\cal H}_{\rm env}$. Such states with reduced energy uncertainty at nonzero excitation energy could be of interest experimentally.
@@
Acknowledgments. I thank S. Gopalakrishnan, D. Huse, R. Jha, S. Sachdev and L. Tagliacozzo for valuable discussions. I am particularly grateful to B. Halperin whose questions led to the investigation of possible non-analytic behavior. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)-217133147/SFB 1073 (Project No. B07). It was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, and at the Kavli Institute for Theoretical Physics (KITP), supported by grants NSF PHY-1748958 and PHY-2309135, whose hospitality is gratefully acknowledged.